Solve the formula for the specified variable.

A= 1/2h(b1 + b2); b
1
Expert's answer
2018-10-05T17:31:09-0400

1) A = h / 2 * (b1 + b2)


2) 2A = h * (b1 + b2)


3) 2A / h = b1 + b2


4) b1 = 2A / h - b2